40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 115 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  11561   Thu Sep 3 00:14:09 2015 ranaConfigurationIOOIMC fast gain change for lock acq

The IMC often was making that scratchy noise when first catching lock and sometimes breaking. Thinking of the crappy crossover sit that EQ showed in his latest plots, I decided that it didn't make sense to acquire lock with an unstable PZT/EOM crossover, so I have changed mcdown to acquire with +13 dB Fast Gain and its much fast now and no longer makes that sound.

I also changed the caput command from 'caput -l' to 'caput -c -l' to see if the async 'wait for callback' feature will insure that the commands get sent. I witnessed the mcdown not actually writing all of its commands once or twice tonight. With the MC Boost left on its never going to lock.

mcdown has been committed to SVN. Please, if you have recently edit mcup and Autolock, commit them to the SVN or else I will delete them and do an svn up.

  11560   Wed Sep 2 23:50:00 2015 ranaUpdatePSLPMC LO dying

Let's order a pair of 35.5 MHz Wenzel for this guy and package like Rich has done for the WB low noise oscillators.

WE're only sending 6 dBm into it now and its using a 13 dBm mixer. Bad for PMC stability.

Also, if anyone has pix of the servo card, please add them to the DCC page for the PMC.

Attachment 1: PMCLO.png
PMCLO.png
  11559   Wed Sep 2 13:44:04 2015 ericqUpdateLSCIR beatnote confusion

From the AFW website about our product, the POBC-64-C-1-7-2-25dB:

port1 slow axis -> port3 slow axis 
port2 slow axis -> port3 fast axis

crying

  11558   Wed Sep 2 01:31:22 2015 ericqUpdateIOOIMC loop shapes

The promised historical comparisons follow. The crossover looks mostly the same as before. There is a new feature in the OLG at 50-60kHz; what could've changed about the EOM path in that time?

 

Attachment 1: imc_olg_trend.pdf
imc_olg_trend.pdf
Attachment 2: imc_xover_trend.pdf
imc_xover_trend.pdf
  11557   Tue Sep 1 17:38:58 2015 ericqUpdateLSCIR beatnote confusion
Quote:

Maybe we just don't understand the splitter/combiners.

After an email from Eric G, I think this is the case.

If you read the text at Thorlabs about Fiber-Based Polarization Beam Combiners/Splitters, it suggests that these things take input beams both aligned to their slow axes, and outputs one field along the slow, and one orthogonal to it on the fast axis. Which is exactly what we don't want for a beat. 

  11556   Tue Sep 1 17:07:06 2015 ericqUpdateLSCIR beatnote confusion

There has been some discussion here and there of using fiber coupled IR beats for ALS. A few weeks ago, and again today with Eric G, I poked around a bit with the fiber box Manasa set up for the FOL scheme. 

Somehow, the IR beatnote is ~1000 times smaller than expected, both with the Thorlabs fiber coupled PD and a fiber coupled NF 1611.

In essence, after the fiber combiner, there is on the order of hundreds of uW each of PSL and AUX X IR light. The output of the fiber from each source looks nice and gaussian. The DC output of the 1611 indicates that it is seeing the right level of light. The green beatnote exists with good SNR at twice the IR beat frequency, so we know that the IR beat isn't some junky modes beating. 

For the 1611, we would expect an RF signal of ~1mW*0.9A/W*700V/A -> .6V / +8dBm. Instead we see ~2mV / -40dBm.

Incidentally, there is some 20mV / -20dBm signal at ~400kHz, presumably from the green PDH modulation at ~200k. 

(The level out of the thorlabs PD is similarly tiny; it doesn't have a DC output though, so we don't know the DC power that the active surface really sees. Not that I expect it to be much different, but the NF just makes it easier to estimate.)

The only things that should be able to cause the beat to be smaller than expected from the power levels are mode matching and polarization matching. All the fibers are single mode, so mode matching should be effectively 100%. Maybe somthing fishy is happening with the polarizations, but they'd have to be really maliciously close to orthogonal to cause this level of mismatch. 

Maybe we just don't understand the splitter/combiners. Mysterious.

  11555   Tue Sep 1 11:56:56 2015 ericqUpdateIOOIMC loop shapes

I took some transfer functions of the IMC loop and crossover, being careful that the PC drive never exceeding 1V during the measurements. 

I then did some algebra to try and back out the individual loop paths, without having to make assumptions/approximations about the loop gain being high enough. This only really works in the region where both the open loop and crossover measurements have coherence. 

It seems to me that the PZT path has pretty low phase margin on its own, but maybe this is ok, since its never really meant to run solo. The EOM path shape is harder to understand.

 

The data I took, and code that made the above plot is attached. This afternoon, I'll post an update comparing the measured OLG and crossover to earlier measurements. 

Attachment 1: IMCshapes_Aug31_2015.pdf
IMCshapes_Aug31_2015.pdf
Attachment 2: IMC_Aug31_2015.zip
  11554   Tue Sep 1 10:36:06 2015 SteveUpdatePEMGurs swapped
Quote:

To help find out if Steve really melted the inside of our precious seismometer, lets hook it up using the handheld seismo wand and see if it produces volts when we shake the ground.

Also, please stop using names like GurA or Gur1 or GurSuzy. We have GurX and GurY because they are at those ends. Anything else is confusing.

I moved Gur A from ETMX to ETMY . Gur B at ETMY was disconnected and its cable  connected to Gur A

It seems that Gur A is alive. I will stop using A and B names after we stop swapping components.

Attachment 1: GurA@ETMY.png
GurA@ETMY.png
  11553   Tue Sep 1 10:26:24 2015 IgnacioUpdateIOOMore MCL Subtractions (Post FF)

Using the training data that was collected during the MISO MCL FF. I decided to look at more MCL subtractions but this time using the accelerometers as Rana suggested.

I first plotted the coherence between MCL and all six accelerometers and the T240-Z seismometer.

For 1 - 5 Hz, based on coherence, I decided to do SISO Wiener filtering with ACC2X and MISO Wiener filtering with ACC2X and ACC1Y. The offline subtractions were as follows (RMS plotted from 0.1 to 10 Hz):

The subtractions above look very much like what you would get offline when using the T240(X,Y) seismometeres during MISO Wiener filtering. But this data was taken with the MISO filters on. This sort of shows the performance deterioration when one does the online subtractions. This is not surprising since the online subtraction performance for the MISO filters, was not too great at 3 Hz. I showed this in some other ELOG but I show it again here for reference:


Anyways, foor 10 - 20 Hz, again based on coherence, I decided to do SISO Wiener filtering with ACC2Z and MISO Wiener filtering with ACC2Z and ACC1Z (RMS plotted from 10 to 20 Hz):

I will try out these subtractions online by today. I'm still debating wether the MISO subtractions shown here are worth the Vectfit shananigans. The SISO subtractions look good enough.

Attachment 4: mclxycoh.png
mclxycoh.png
  11552   Tue Sep 1 06:58:11 2015 IgnacioUpdateWienerFilteringMCL FF => WFS1 and WFS2 FF => ARMS FF

I took some training data during Sunday night/Monday morning while the MCL MISO FF was turned on. We wanted to see how much residual noise was left in the WFS1/WFS2 YAW and PITCH signals. 

The offline subtractions that can be achieved are:

For WFS1

 

For WFS2

I need to download data for these signals while the MCL FF is off in order to measure how much subtraction was achived indirectly with the MCL FF. In a previous elog:11472, I showed some offline subtractions for the WFS1 YAW and PITCH before any online FF was implemented either by me or Jessica. From the plots of that eLOG, one can clearly see that the YAW1 signal is clearly unchanged in the sense of how much seismic noise was mitigated indirectly torugh MCL. 

Koji has implemented the FF paths (thank you based Koji) necessary for these subtractions to be implemented. The thing to figure out now is where we want to actually actuate and to measure the corresponding transfer functions. I will try to have either Koji or Eric help me measure some of these transfer functions.

Finally, I looked at the ARMS and see what residual seismic noise can be subtracted

 

I'm not too concerned about noise in the arms as if the WFS subtractions turn out to be promising then I expect for some of the arms seismic noise to go down a bit further. We also don't need to measure an actuator transfer function for arm subtractions, give that its essentially flat at low frequencies, (less than 50 Hz).

 

  11551   Tue Sep 1 02:44:44 2015 KojiSummaryCDSc1oaf, c1mcs modified for the IMC angular FF

[Koji, Ignacio]

In order to allow us to work on the IMC angular FF, we made the signal paths from PEM to MC SUSs.
In fact, there already were the paths from c1pem to c1oaf. So, the new paths were made from c1oaf to c1mcs. (Attachment 1~3)

After some debugging those two models started running. The additional cost of the processing time is insignificant.
FB was restarted to accomodate the change.

Once the modification of the models was completed, the OAF screens were modified. It seemed that the Kissel button
for the output matrix haven't been updated for the PRM ASC implementation. This was fixed as the button was updated this time.
In addition, the button for the FM matrix was also made and pasted.

 

Attachment 1: c1oaf_screenshot1.png
c1oaf_screenshot1.png
Attachment 2: c1oaf_screenshot2.png
c1oaf_screenshot2.png
Attachment 3: c1mcs_screenshot.png
c1mcs_screenshot.png
Attachment 4: OAF_MEDM1.png
OAF_MEDM1.png
Attachment 5: OAF_MEDM2.png
OAF_MEDM2.png
  11550   Mon Aug 31 14:15:23 2015 IgnacioUpdateIOOMeasured the MC_F whitening poles/zeroes

I measured the 15 Hz zero and the 150 Hz pole for the whitening filter channels of the Generic Pentek board in the IOO rack. The table below gives these zero/pole pairs for each of the 8 channels of the board.

channel zero [Hz] pole [Hz] Chan
1 15.02 151.05 C1:ASC-POP_QPD_YAW
2 15.09 150.29 C1:ASC-POP_QPD_PIT
3 14.98 150.69 C1:ASC-POP_QPD_SUM
4 14.91 147.65 C1:ALS-TRX
5 15.03 151.19 C1:ALS-TRY
6 15.01 150.51 ---
7 14.95 150.50 C1:IOO-MC_L
8 15.03 150.93 C1:IOO-MC_F

Here is a plot of one of the measured transfer functions,

and the measured data is attached here: Data.zip


EQ: I've added the current channels going through this board. 

More importantly, I found that the jumpers on channel one (QPD X) were set to no whitening, in contrast to all other channels. Thus, the POP QPD YAW signals we've been using for who knows how long have been distorted by dewhitening. This has now been fixed. 

Hence, the current state of this board is that the first whitening stage is disabled for all channels and the second stage is engaged, with the above parameters. 

Attachment 1: Data.zip
  11549   Mon Aug 31 09:36:05 2015 IgnacioUpdateIOOMISO Wiener Filtering of MCL

MISO Wiener filters for MCL kept the mode cleaner locked for a good 8+ hours.

  11548   Mon Aug 31 07:49:11 2015 ericqUpdateIOOMC2 -> MCL Actuator TF
Quote:

I think what happened here is you forgot to undo the MC_F whitening filter which is the Generic Pentek Interface board next to the MC servo board. I suggest you guys measure this on Monday so you can correctly estimate the MC length noise. And then perhaps undo the whitening in the anti-whitening filter of this filter bank so that the signal which is recorded is in units of kHz.

This should allow your online subtraction filter to be more correct: roughly speaking, the phase shift below a pole or zero is going to be 45*(f/fp) deg. Since we expect there to be 2 zeros at 15 Hz, it would be 9 deg phase shift at 1.5 Hz and limit the subtraction to ~80%.

While it is true that the whitening filter was incorrectly handled, I don't think this should change the subtraction performance since the MC_L data used for the Wiener filter training was also taken without undoing the whitening filter.

  11547   Sun Aug 30 23:47:02 2015 IgnacioUpdateIOOMISO Wiener Filtering of MCL

I decided to give MISO Wiener filtering a try again. This time around I managed to get working filters. The overall performance of these MISO filters is much better than the SISO I constructed on elog:11541 .

The procedure I used to develope the SISO filters did not work well for the construction of these MISO filters. I found a way, even more systematic than what I had before to work around Vectfit's annoyances and get the filters in working condition. I'll explain it in another eLOG post.

Anyways, here are the MISO filters for MCL using the T240-X and T240-Y as witnesses:

 Now the theoretical offline prediction:

 

 

The online subtractions for MCL, YARM and XARM. I show the SISO subtraction for reference.

 And the subtraction performance:

  11546   Sun Aug 30 13:55:09 2015 IgnacioUpdateIOOSummary pages MCF

The summary pages show the effect of the MCL FF on MCF (left Aug 26, right Aug 30):

 

I'm not too sure what you meant by plotting the X & Y arm control signals with only the MCL filter ON/OFF. Do you mean plotting the control signals with ONLY the T-240Y MCL FF filter on/off? The one that reduced noise at 1Hz?

 

 

  11545   Sun Aug 30 13:31:48 2015 ranaUpdateIOOMCL Wiener Feedforward Final Results

I'm not totally sure, but by eyeball, this seems like the best online MCL reduction we've ever had. Nice work.

The 3 Hz performance is the same as usual, but we've never had such good 1 Hz reduction in the online subtraction.

I would like to see a plot of the X & Y arm control signals with only the MCL filter ON/OFF. This would tell us how much of the arm signals were truly frequency noise.

  11544   Sun Aug 30 12:20:08 2015 ericqUpdateIOOMCL Wiener Feedforward Final Results
Quote:

Big thumbnails? Could it have been this? elog:11498.

Ignacio is correct; I forgot to shrink the value back down after testing the PDF thumbnails. Default thumbnail size is now back to 600px. 

  11543   Sun Aug 30 10:57:29 2015 IgnacioUpdateIOOMCL Wiener Feedforward Final Results

Big thumbnails? Could it have been this? elog:11498.

Anyways, I fixed the plots and plotted an RMS that can actaully be read in my original eLOG. I'll see what can be done with the MC1 and MC2 Wilcoxon (z-channel) for online subtractions. 

  11542   Sun Aug 30 00:03:13 2015 ranaUpdateIOOMCL Wiener Feedforward Final Results

Somehow it seems like the ELOG makes all of the thumbnails way too big by default. Did we get some sneaky upgrade recently?

I would only plot your results below 50 Hz. We don't care about the RMS at high frequencies and it can make the RMS misleading.

We definitely need to include one vertical Wilconox at each MC chamber so that it can subtract all of that junk at 10-20 Hz.

  11541   Sat Aug 29 04:53:24 2015 IgnacioUpdateIOOMCL Wiener Feedforward Final Results

After fighting relentlessly with the mode cleaner, I believe I have achieved final results

I have mostly been focusing on Wiener filtering MCL with a SISO Wiener filter for a reason, simplicity. This simplicity allowed me to understand the dificulties of getting a filter to work on the online system properly and to develope a systematic way of making this online Wiener filters. The next logical step after achieving my final SISO Wiener filter using the T240-X seismometer as witness for MCL (see elog:11535) and learning how to produce good conditioned Wiener filters was to give MISO Wiener filtering of MCL a try. 

I tried performing some MISO filtering on MCL using the T240-X and T240-Y as witnesses but the procedure that I used to develope the Wiener filters did not work as well here. I made the decision to ditch it and use some of the training data I saved when the SISO (T240-X) filter was runing overnight to develope another SISO Wiener filter for MCL but this time using T240-Y as witness. I will compare how much more we gain when doing MISO Wiener filtering compared to just a bunch of SISO filtering in series, maybe a lot, or little.

I left both filters running overnight in order to get trainining data for arm and WFS yaw and pitch subtractions.

The SISO filters for MCL are shown below:

The theoretical FIR and IIR subtractions using the above filters:

 

Running the filters on the online system gave the following subtractions for MCL and YARM:

 

Comparing the subtractions using only the T240-X filter versus the T240-X and T240-Y:

 

 

  11540   Fri Aug 28 20:20:26 2015 ranaUpdatePEMGur interface box cable

To help find out if Steve really melted the inside of our precious seismometer, lets hook it up using the handheld seismo wand and see if it produces volts when we shake the ground.

Also, please stop using names like GurA or Gur1 or GurSuzy. We have GurX and GurY because they are at those ends. Anything else is confusing.

  11539   Fri Aug 28 20:15:49 2015 ranaUpdateIOOMC2 -> MCL Actuator TF
qQuote:

I made a measurement of the MC2 actuator transfer function by injecting noise from 1-100Hz into LSC_MC2_EXC for about 15 minutes, then estimating the TF from MC2_OUT to IOO_MC_L with CSD/PSD. The inverse of this TF will be applied to their Wiener target data to give us the direct subtration filter we want.

I think what happened here is you forgot to undo the MC_F whitening filter which is the Generic Pentek Interface board next to the MC servo board. I suggest you guys measure this on Monday so you can correctly estimate the MC length noise. And then perhaps undo the whitening in the anti-whitening filter of this filter bank so that the signal which is recorded is in units of kHz.

This should allow your online subtraction filter to be more correct: roughly speaking, the phase shift below a pole or zero is going to be 45*(f/fp) deg. Since we expect there to be 2 zeros at 15 Hz, it would be 9 deg phase shift at 1.5 Hz and limit the subtraction to ~80%.

  11538   Fri Aug 28 19:05:53 2015 ranaUpdateIOOIMC Tweak

Well, green looks better than blue, but it makes the PCDRIVE go high, which means its starting to saturate the EOM drive. So we can't just maximize the phase margin in the PZT/EOM crossover. We have to take into account the EOM drive spectrum and its RMS.

Also, your gain bump seems suspicious. See my TF measurements of the crossover in December. Maybe you were saturating the EOM in your TF ?

Lets find out what's happening with FSS servos over in Bridge and then modify ours to be less unstable.

  11537   Fri Aug 28 17:04:12 2015 SteveUpdatePEMGur interface box cable
Ara! Onara suru tsu-mori datta keh-do, un-chi ga de-chatta...
Quote:

 

Quote:

As reported previously, the transfer functions of the channels look fine. (i.e. All channels almost identical.)
I checked the chain from the unit input to the DAQ BNC connectors. They were all OK.

Today I have been checking the signals on the unit with the long DB37 cables connected.
I could not see anything on the Gur2 channels on the board. I looked at the DB37 for Gur2 and felt something is wrong.

I opened the housing of the cable and realized that all the pins are not fully inserted.
The wires were crimped improperly and prevents them to be fully inserted.

=> We need to redo the crimping to insert them.
=> We need to check the other side too.

I'm making a new long cable. Both connector ends of this X arm long cable were terrible. It was removed from the cable tray yesterday.

The Y arm Guralp is running fine. The interface box is open in front of 1X1 on a cart and it is alive!    Please be aware of it !

The new cable was made this way:

Pins were located with ohm meter for locations and both ends were cut off.The Belden 1424A cable than was soldered to DB37 and Guralp circular connector " IKPT06F16-26S-ND "

This cable will connect the ADC interface box CCD# DO 60506 to Guralp seismometer at the sout end.

The Guralp's each 3 axies will be connected through a twisted pair to the differential input op amp

Gur  ouput  vertical axis + -  on circular connector   A, B  to DB37 pin 1 &20

Gur output        N/S axis + -  on circular connector   C, D to DB37 pin 2 & 21

Gur output       E/W axis + -  on circular connector   E, F to DB37 pin 3 & 22

Power +12 VDC  from DB37 pin 29 to Gur circular connector pin c (lower case)

Power -12 VDC  from DB37 pin 24 to Gur circular connector pin  M Note: this connection was absent at the first test of this cable!

Ground              from DB37 pin 10 to Gur circular connector pin b (lower case)

Summery: I may destored the opamp  at the Guralp A at the south end.

 

 

Attachment 1: fixedNewLongCable.png
fixedNewLongCable.png
  11536   Fri Aug 28 02:20:35 2015 IgnacioUpdateLSCPRFPMI and MCL FF

A day late but here it is.

Eric and I turned on my SISO MCL Wiener filter elog:11535 during his PRFPMI 40min lock. We looked at the CARM_IN and CARM_OUT signals during the lock and with the MCL FF on/off. Here is the spectra:

  11535   Fri Aug 28 00:59:55 2015 IgnacioUpdateIOOFinal SISO FF Wiener Filter for MCL

This is my final SISO Wiener filter for MCL that uses the T240-X seismo as its witness.

The main difference between this filter and the one on elog:11532 is the actual 1/f rolloff this filter pocesses. My last filter had a pair of complex zeroes at 2kHz, that gave the filter some unusual behavior at high frequencies, thanks Vectfit. This filter has 10 poles and 8 zeroes, something Vectfit doesn't allow for and needs to be done manually.

The nice thing about this filter is the fact that Eric and I turned this filter on during his 40 min PRFPMI lock last night, Spectra for this is coming soon.

This filter lives on the static Wiener path on the OAF machine, MCL to MC2, filter bank 7.

Anyways, the usual plots are shown below. 

 

Filter:

T240-X (SISO)

 

Training data + Predicted FIR and IIR subtraction:

 

Online subtraction results:(High freq. stuff shown for noise injection evaluation of the filter)

MCL

YARM

Subtraction performace:

  11534   Thu Aug 27 04:23:04 2015 ericqUpdateLSCPRFPMI is back

Got to a 40 minute lock tonight. All other locks broke because of me poking something. 

I redid some sensing excitations, right after carefully measuring the CARM OLG at its excitation frequency, so I can get at the open loop PD response. 

I also used a MCL feedforward filter of Ignacio's which did not inject any observable noise into the CARM error signal during PRFPMI lock. He will make some elog about this. 

  11533   Thu Aug 27 02:09:14 2015 ericqUpdateLSCAUX X Laser Current Changed

I spent some time tonight chasing down the cause of huge RIN in the X green PDH transmitted light, which I had started seeing on Monday. This was preventing robust locking, since the ALS sensing noise was ~10x worse above 50Hz, thus making the AO transition much flakier (though, impressively, not impossible!)

I went down to the X end, and found that turning the laser diode current down by 0.1A (from 2.0 to 1.9) smoothed things out completely. Unfortunately, this causes the power to drop, from GTRX of 0.45 to 0.3, but the ALSX sensitivity is unchanged, as compared with the recenent "out of loop" template. 

This also seems to have changed the temperatures of the good modes, as no beat was evident at the previously good temperature. Beats were found at +5400 and +10500 counts on the slow servo offset slider; I suspect the third lies around the edge of the DAC range which is why I couldn't uncover it. In any case, I've parked it at 10500 for now, and will continue locking; nailing it down more precisely and offloading the slider offset to the laser controller will happen during daytime work...

  11532   Thu Aug 27 01:41:41 2015 IgnacioUpdateIOOTriply Improved SISO (T240-X) FF of MCL

Earlier today I constructed yet another SISO filter for MCL. The one thing that stands out about this filter is its strong roll off wink. This prevents high frequency noise injection into YARM. The caviat, filter performance suffered broken heart quite a bit, but there is subtraction going on.

I have realized that Vectfit lacks the ability of constraining the fits it produces, (AC coupling, rolloff, etc) even with very nitpicky weighting. So the way I used vectfit to produce this filter will be explained in a future eLOG, I think it might be promising. 

Anyways, the usual plots are shown below. 

 

Filter:

T240-X (SISO)

 

 

Training data + Predicted FIR and IIR subtraction:

 

Online subtraction results:(High freq. stuff shown for noise injection evaluation of the filter)

MCL
 
YARM
 
 
 

Subtraction performace:

 

  11531   Tue Aug 25 16:39:06 2015 SteveUpdatesafetysafety training

Alessandra Marrocchesi received 40m specific basic safety training yesteday.

  11530   Tue Aug 25 16:33:31 2015 Ignacio, SteveConfigurationPEMSeismometer enclosure copper foil progress

Steve ordered about two weeks ago a roll of 0.5 mm thick copper foil to be used for the inside of the seismometer cans. The foil was then waterjet cut by someone in Burbank to the right dimensions (in two pieces, a side and a bottom for each of the three cans).

Today, we glued the copper foil (sides only) inside the three seismometer cans. We used HYSOL EE4215/HD3561(Data Sheet) as our glue. It is a "high impact, low viscocity, room temperature cure casting" that offers "improved thermal conductivity and increased resistance to heat and thermal shock." According to Steve, this is used in electronic boards to glue components when you want it to be thermal conductive.

We are going to finish this off tomorrow by gluing the bottom foil to the cans. The step after this involves soldering the side to the bottow and where the side connects. We have realized that the thermal conductivity of the solder that we are using is only ~50. This is 8 times smaller than that of copper and wil probably limit how good a temperature gradient we will have.

Some action shots,

  11529   Tue Aug 25 16:09:54 2015 ericqUpdateIOOIMC Tweak

A little more information about the IMC loop tweak...

I increased the overall IMC loop gain by 4dB, and decreased the FAST gain (which determines the PZT/EOM crossover) by 3dB. This changed the AO transfer function from the blue trace to the green trace in the first plot. This changed the CARM loop open loop TF shape from the unfortunate blue shape to the more pleasing green shape in the second plot. The red trace is the addition of one super boost. 

 

Oddly, these transfer functions look a bit different than what I measured in March (ELOG 11167), which itself differed from the shaping done December of 2014 (ELOG 10841). 


I haven't yet attempted any 1F handoff of the PRMI since relocking, but back when Jenne and I did so in April, the lock was definitely less stable. My suspicion is that we may need more CARM supression; we never computed the loop gain requirement that ensures that the residual CARM fluctuations witnessed by, say, REFL55 are small enough to use as a reliable PRMI sensor.

I should be able to come up with this with data from last night. 

Attachment 1: imcTweak.pdf
imcTweak.pdf
Attachment 2: CARM_TF.pdf
CARM_TF.pdf
  11528   Tue Aug 25 04:15:51 2015 ericqUpdateLSCPRFPMI is back

More PRFPMI locks tonight. Right now, it's been locked for 22+ minutes, though with the PRMI still on 3F signals. I think the MC2/AO crossover needs some reshaping; there's a whole bunch of noise injected into CARM around 600 Hz, which is where the two paths differ by 180deg. (Addendum: broke lock at ~27 minutes, 4:16AM)

For most of this lock, sensing matrix excitations have been running for daytime analysis. 

The nominal IMC loop gain / EOM crossover were making the AO path very marginal. I've adjusted the nominal settings and autolocker scripts. 

There was some weird behavior of X green PDH earlier... Broadband RIN seen in ALS-TRX, coherent with the DC output of the beat PD, so really on the light. I fiddled with the end setup, and it mostly went away, though I didn't intentionally change anything. Disconcerting. 

  11527   Mon Aug 24 16:46:49 2015 ericqUpdateSUSETMY Oplev laser power is falling

Repeated for all optics, ETMY seems like the only one sharply dropping for now (PRM is all over the place and hard to gauge, since we often leave it partially- or mis-aligned):

 


Hacky bits:

Bash:

find /opt/rtcds/caltech/c1/burt/autoburt/snapshots/2015 -wholename "*00:0*"
| xargs ack --nogroup "OL_SUM_OUT16 1"
| grep -v 'SUS-MC'
| sed -e 's/.*2015/2015/g' -e 's/\/c1.*C1:SUS-/, /g' -e 's/_OL.*\([0-9]\..*$\)/, \1/g' -e 's/\//-/g'
| sort | uniq > allOL.txt

qontrols@pianosa|~ > head allOL.txt 
2015-Apr-10-00:07, BS, 1.146766113281250e+03
2015-Apr-10-00:07, ETMX, 1.597261328125000e+04
2015-Apr-10-00:07, ETMY, 4.331762207031250e+03
2015-Apr-10-00:07, ITMX, 6.488521484375000e+03
2015-Apr-10-00:07, ITMY, 1.387590234375000e+04
2015-Apr-10-00:07, PRM, 8.352053833007812e+02
2015-Apr-10-00:07, SRM, 6.099560928344727e+01
2015-Apr-1-00:07, BS, 1.180478149414062e+03
2015-Apr-1-00:07, ETMX, 1.584842480468750e+04

Python:

 

olsum = pd.read_csv('allOL.txt',parse_dates=True, names=['Date','Optic','Sum'])
olsum['Date'] = pd.to_datetime(olsum['Date']) # Automatic parsing didn't work for some reason
olpivot = olsum.pivot(index='Date',columns='Optic',values='Sum')
olpivot.plot()
Attachment 1: OLsum_trend.pdf
OLsum_trend.pdf
  11526   Mon Aug 24 16:10:07 2015 ericqUpdateSUSETMY Oplev laser power is falling

Today I noticed the box around the ETMY oplev sum flashing red, as it dipped below 1k. I don't recall seeing this recently, so I wanted to look up the history.

However, we've been having trouble with our minute (and longer) trend data, so I had to hack it out a bit... Here is the unfortunate result:

I think we can be fairly confident that this is not due to alignment drifts, we generally keep the QPD reasonably well centered. I also recentered it today, and the counts remained at ~1k. 


Details of the hack that got me this data:

I ended up looking at the BURT snapshots from every night at midnight, which report a number for ETMY_OL_SUM_OUT16, and making a text file with dates and values with the following BASH spaghetti:

find /opt/rtcds/caltech/c1/burt/autoburt/snapshots/2015 -wholename "*00*/*scy*" |xargs ack --nogroup "ETMY_OL_SUM_OUT16 1" |  sed -e 's/.*2015/2015/g' -e 's/\/c1.*\([0-9]\..*$\)/, \1/g' -e 's/\//-/g'  > ETMYsum.txt

This produces a file full of unsorted lines like: 2015-Aug-23-00:07, 1.106459228515625e+03

The python package pandas is good at parsing dates and automatically plotting time series: 

olsum = pandas.read_csv('ETMYsum.txt', index_col=0, parse_dates=True)
olsum.plot()
Attachment 1: ETMYsum_trend.pdf
ETMYsum_trend.pdf
  11525   Mon Aug 24 14:05:47 2015 ericqUpdateSUSEricG Investigating L2A

This afternoon, I showed Eric Gustafson some of the basics of making swept sine measurements with DTT. We turned off the f2a filters and oplev damping on the BS and made a cursory measurement of the transfer function from position drive to the oplev signals. 

He will be in the lab periodically to continue this line of investigations. 

  11524   Sat Aug 22 15:48:32 2015 KojiSummaryLSCArm locking recovery

As per Ignacio's request, I restored the arm locking.

- MC WFS relief

- Slow DC restored to ~0V

- Turned off DARM/CARM

- XARM/YARM turned on

- XARM/YARM ASS& Offset offloading

  11523   Fri Aug 21 17:15:13 2015 SteveUpdatePEMGur interface box

 

Quote:

As reported previously, the transfer functions of the channels look fine. (i.e. All channels almost identical.)
I checked the chain from the unit input to the DAQ BNC connectors. They were all OK.

Today I have been checking the signals on the unit with the long DB37 cables connected.
I could not see anything on the Gur2 channels on the board. I looked at the DB37 for Gur2 and felt something is wrong.

I opened the housing of the cable and realized that all the pins are not fully inserted.
The wires were crimped improperly and prevents them to be fully inserted.

=> We need to redo the crimping to insert them.
=> We need to check the other side too.

I'm making a new long cable. Both connector ends of this X arm long cable were terrible. It was removed from the cable tray yesterday.

The Y arm Guralp is running fine. The interface box is open in front of 1X1 on a cart and it is alive!    Please be aware of it !

  11522   Fri Aug 21 08:38:37 2015 SteveUpdateIOOPMC locked

PMC needed to be locked manually.cheeky

Attachment 1: PMClocked.png
PMClocked.png
  11521   Thu Aug 20 18:08:28 2015 IgnacioFrogs40m upgradingFatality. Something broke.

So I made coffee at 1547 and was astonished to find the above. Its a sad, very sad day.

At first I thought that something (a gravity wave?) or someone, accidentally hit the thing and it fell and broke. But Koji told me that the janitor was cleaning around the thing and it did indeed fell accidentally.

  11520   Thu Aug 20 11:31:55 2015 KojiUpdatePEMGur interface box

As reported previously, the transfer functions of the channels look fine. (i.e. All channels almost identical.)
I checked the chain from the unit input to the DAQ BNC connectors. They were all OK.

Today I have been checking the signals on the unit with the long DB37 cables connected.
I could not see anything on the Gur2 channels on the board. I looked at the DB37 for Gur2 and felt something is wrong.

I opened the housing of the cable and realized that all the pins are not fully inserted.
The wires were crimped improperly and prevents them to be fully inserted.

=> We need to redo the crimping to insert them.
=> We need to check the other side too.

Attachment 1: IMG_1958.JPG
IMG_1958.JPG
  11519   Thu Aug 20 11:09:10 2015 ranaUpdateIOOsome points about seismic FF
  • When plotting the subtraction performance, we mainly care about the 0.5 - 10 Hz band, so we care about the RMS in this band. Don't integrate over the whole band.
  • When calculating the Wiener filter, you must use the pre-weighting so as to not let the Wiener residual be dominated by the out of band signals. We don't want the filter to try to do anything outside of the 0.5 - 10 Hz band.
  • Somehow, we want to assign a penalty for the filter to have high frequency gain. We do NOT want to slap on an ad-hoc low pass filter. The point of the Wiener filtering is to make the optimum.
  • What is the reason for the poor filter performance from 0.5 - 2 Hz ? If we use the frequency domain (Dmass) subtraction technique, we can do better, so there's some inefficiency in this process.
  • we're getting too much of the 3 Hz stack mode coupling into MCL. I think this means that our damping filters should be using RG around the suspension eigenmodes rather than just simple velocity damping. We had this years ago, but it caused some weird interaction with the angular loops...to be puzzled out.
  11518   Thu Aug 20 02:31:09 2015 ericqUpdateLSCPRFPMI is back

PRFPMI locking has been revived.

I've had 6 5min+ locks so far; arm powers usually hit ~125 for a recycling gain of about 7; visibility is about 75%

The locking script takes a little under 4 minutes to take you from POX/POY lock to PRFPMI if you don't have to stop and adjust anything.

At Koji's suggestion, I used digital REFL11 instead of CM_SLOW, which got me to a semistable lock with some RF, at which time I could check the CM_SLOW situtation. It seemed like the whitening Binary IO switch got out of sync with the digital FM status somehow... 

I've been making the neccesary changes to the carm_cm_up script. I also added a small script which uses the magnitude of the I and Q signals to set the phase tracker gain automatically based on some algebra Koji posted in an ELOG some years ago. 

The RF transition seems much smoother now, most likely due to the improved PRC and ALS stability. In fact, it is possible to hold at arm powers of >100 solely on the digital servos; I don't think we were able to do this before until the AO had kicked in. 

Right now I'm losing lock when trying to engage the CARM super boost. I also haven't switched the PRMI over to 1F signals yet. Would be good to hook the SR785 back up for a loop TF, but I'll stop here for tonight since our SURFs are presenting bright and early tomorrow morning. 

Attachment 1: lock.pdf
lock.pdf
  11517   Wed Aug 19 07:58:25 2015 SteveUpdatePEMGur interface box

Koji and Steve,

We took transferfunctions of each channel yesterday. They were identical ?. I will check the cables from ADC to DAQ next.

Attachment 1: GurADCbox.jpg
GurADCbox.jpg
  11516   Wed Aug 19 01:45:10 2015 IgnacioUpdateIOODoubly Improved SISO (T240-X) FF of MCL

Today I tried and doubly-improved SISO FF filter on MCL. This filter has a stronger rolloff than the previous SISO filters I have tried. The rolloff most definelty helped towards reducing the ammount of noise being injected into YARM. Below is the usual stuff:

 

Filter:

T240-X (SISO)

 

 

Training data + Predicted FIR and IIR subtraction:

 

Online subtraction results:

MCL
YARM
MCL TRANS
 
 
 

Subtraction performace:

  11515   Wed Aug 19 00:55:35 2015 IgnacioUpdateLSCLSC-YARM-EXC to LSC-YARM-IN1 TF measurement + error analysis

Yesterday, Rana, Jessica and I measured the Transfer function from LSC-YARM-EXC to LSC-YARM-IN1. 

The plot below shows the magnitude and the phase of the measured transfer function. It also shows the normalized standard error in the estimated transfer function magnitude; the same quantity can be applied to the phase, only in this case it is interpreted as its standard deviation (not normalized). It is given by

 \frac{[1-\gamma_{xy}^2(f)]^{1/2}}{|\gamma_{xy}(f)|\sqrt{2n_{d}}}

where \gamma_{xy}^2(f) is the ordinary coherence function and n_{d} is the number of averages used at each point of the estimate, in the case here we used 9 averages. This quantity is of interest to us in order to understand how the accuracy of transfer function measurement affects the ammount of subtraction that can be achieved online.

 

Since this transfer function is flat from 1-10 Hz (out of phase by 180 deg), this means that we can apply our IIR wiener filters direclty into YARM without taking into account the TF by prefiltering our witnesses with it. Of course this is not the case if we care about subtractions at frequencies higher than 10 Hz, but since we are dealing with seismic noise this is not a concern.

The coherence for this transfer function measurement is shown below,

  11514   Tue Aug 18 11:16:17 2015 SteveUpdatePEMGur interface box is wonky

The Guralp ADC  interface box  D060506 is ready for inspection. It is in front of 1X1 with open top and running.cool

Obviously c7 as 1 miroF cap is missing.blush

Quote:

Let's dismantle the I/F unit from the rack and connect the cable with the lid open.
We need to trace the signal.

 

Attachment 1: IMG_0009.JPG
IMG_0009.JPG
  11513   Tue Aug 18 03:56:09 2015 ericqUpdateLSClocking efforts

Now that the updated ALS is stable, and the PRC angular FF is revived, I've been working on relocking PRFPMI. While the RMS arm fluctuations are surely smaller than they used to be, there is no noticible difference to the ears when buzzing around resonance, but this doesn't really mean much. 

Frustratingly, I am not able to stably blend in any RF CARM error signal into the slow length control path (i.e. CARM_B). Bringing AS55 Q into DARM with the 20:0 integrator is working fine, but we really need to supress CARM to get anywhere. I'm not sure why this isn't working; poking around into the settings that were used when we were regularly locking didn't turn up any differences as far as I could tell. Investigations continue... 

Some minor changes to the locking script were made, to account for the increased ALS displacement sensitivity from the longer delay line. 


Since the ALS is now in a fairly stable state, I've updated the calibrated PSD template at /users/Templates/ALS/ALS_outOfLoop_Ref.xml, and added some coherence plots for some commonly coupled quantities (beat signal amplitude, IR error signal, green PDH error signal and green transmission). 

Attachment 1: newALSref.pdf
newALSref.pdf
Attachment 2: xCoh.pdf
xCoh.pdf
Attachment 3: yCoh.pdf
yCoh.pdf
  11512   Mon Aug 17 17:48:12 2015 KojiUpdatePEMWasps obliterated maybe...

We found the same wasp in the 40m. Megan found it walking behind Steve desk!

ELOG V3.1.3-